A 15-A current flows device that has a resistance of 0.15 ohm's. Compute the rate of heat production by the device


1
Expert's answer
2021-04-22T07:35:39-0400

The rate of heat production is "P = I^2 R = (15 A)^2 \cdot 0.15 \Omega = 33.75 W".